OpenStudy (anonymous):

Find the value of y when x equals -2 in this equation. -3x + 3y = -12

OpenStudy (anonymous):

y=-6

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

Your welcome(:

OpenStudy (anonymous):

u want it explained???

OpenStudy (anonymous):

yeah because i have no idea how to do this crapp!

OpenStudy (anonymous):

figured that i cant handle just the answer

OpenStudy (anonymous):

so first of all u do -3x + 3y = -12////-3^-2 + 3y = -12

OpenStudy (anonymous):

okay!!

OpenStudy (anonymous):

so can u figure out the first part you have to solve?

OpenStudy (anonymous):

the one after the ///////

OpenStudy (anonymous):

ok good so first solve -3^-2

OpenStudy (anonymous):

9

OpenStudy (anonymous):

srry i back and forth with something

OpenStudy (anonymous):

k

OpenStudy (anonymous):

and its accually 6 :P

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so 6+3y=-12

OpenStudy (anonymous):

do you know what we do now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so we subtract 6 to both sides of the equation

OpenStudy (anonymous):

so 6+3y=-12 -6 -6 3y=-18

OpenStudy (anonymous):

do u think u can finish????